PLEASEEEEEE HELPPPPP
If paul takes 10hours to paint a house , and wiliam takes 5 hours to paint the same house. how much time it will take them to paint the house if they work together ???

Answer by stanbon(75887)

If paul takes 10hours to paint a house , and wiliam takes 5 hours to paint the same house. how much time it will take them to paint the house if they work together ???
-------------
Paul's rate: 1/10 job/hr
---
William's rate: 1/5 job/hr
--------------
Together rate: 1/x job/hr
==============================
Equation:
rate + rate = together rate
1/10 + 1/5 = 1/x
---
Multiply thru by 10x:
x + 2x = 10
3x = 10
x = 3 1/3 hrs = 3 hr 20 min (time to do the job together)
====================
Cheers,
Stan H.